Unit mass of certain fluid is contained in a




cylinder at an initial pressure of 20 bar. The fluid




is allowed to expand reversibly behind a piston




according to a law Pv2=constant until the volume




is doubled. The fluid is then cooled reversibly at




constant pressure until the piston regains its




original position; heat is then supplied reversibly




with the piston firmly locked in position until the




pressure rises to the original value of 20 bar.




Calculate the net work done by the fluid, for an




initial volume of 0.05m

1
Expert's answer
2022-01-05T05:21:02-0500

"V_1=0.05m^3"

"V_2=2V_1=2*0.05=0.1m^3"

"P_2=P_1(\frac{V_1}{V_2})^n" , Where n=2

"P_2=20(\frac{1}{2})^2=5bar"

"P_3=P_2=5bar"

"V_3=V_1=0.05m^3"

"W=W_{1-2}+W_{2-3}+W_{3-1}"

"=\frac{P_1V_1-P_2V_2}{n-1}+P_2(V_3-V_2)+0"

"=\frac{20*10^2*0.05-5*10^2*0.1}{2-1}+5*10^2(0.05-0.1)"

=25kJ